A helicopter spraying fertilizer over a field can fly 0.4 mile downwind in the same time as it can fly 0.3 mile upwind. Find the speed of the wind if the helicopter travels 35 mph in still air when dusting crops.


____________mph

Answer by mananth(16949)

Helicopter speed = 35 mph
wind speed given = x mph

speed against wind 35 - x
speed with wind 35 + x

Distance against wind 0.3 miles
Distance with wind 0.4 miles

t=d/r time against = time with wind -
0.3 / ( 35 - x )= 0.4 / ( 35 + x )
0.3 ( 35 + x )= 0.4 ( 35 -x )
10.5 + 0.3 x = 14 - -0.4 x
0.4 x 0.3 x 14 + -10.5
0.7 x = 3.5
/ 0.7
x = 5 mph wind speed